OpenStudy (anonymous):

|x-9| = 3 what is x????

OpenStudy (anonymous):

its 6? or 12?

OpenStudy (anonymous):

|x-9| = 3 so since you have the absolute value sign then you have two possible answers x-9 = 3 or x-9 = -3 so for x-9 = 3 x-9 = 3 x = 9+3 = 12 OR x-9 = -3 x = -3-9 x = -12

OpenStudy (anonymous):

Ahh ok! Thanks :D

OpenStudy (anonymous):

haha sorry, did that wrong

OpenStudy (anonymous):

oh ok its fine